At what price should a dealer mark an aquarium in order that the selling price may be $980 after allowing a discount of 15% on the marked price???
Please help...

Answer by Theo(3464)

dealer wants the selling price to be $980.
this price should include a discount of 15% on the marked price.
this means that the marked price minus 15% should be equal to $980.
let x equal the marked price.
then x - .15 * x = 980.
factor out the x to get (1-.15)*x = 980.
simplify to get .85*x = 980.
divide both sides of the equation by .85 to get x = 980/.85
simplify to get x = 1152.941176.
that's what the marked price should be.
the aquarium is marked as costing 1152.941176.
with a discount of 15%, the selling price will be 1152.941176 minus .15 * 1152.941176, which becomes 980.